# Shift Happens: Mixture of Experts based Continual Adaptation in Federated Learning

### v1

@@ 7. Results and Discussion
We present the numerical results for Adaptation Time, Accuracy Drop, and Max Accuracy in Tables 1 and 2. To visualize convergence behavior, we plot Accuracy (%) over communication rounds in Figures 3(a) through 4(b). Figures 5(a) through 6(b) show the peak accuracy achieved by each method in every window. Finally, to illustrate how ShiftEx dynamically adapts to distributional changes, we visualize expert assignment distributions across windows in Figures 7(a) through 8(b). These plots show how ShiftEx begins with a single expert in W0 and progressively creates and trains new experts as shifts are detected, enabling specialized adaptation to new data regimes.
ShiftEx reduces adaptation time under distribution shifts. Across all datasets, ShiftEx achieves significantly faster adaptation than all baselines. On the large-scale FMoW dataset (Table 1), where most baselines fail to recover within 51 rounds, ShiftEx consistently recovers within 1–13 rounds across all windows. In TinyImagenet-C (Table 2), recovery time drops to 27–29 rounds, whereas other methods plateau with no recovery even after 40 rounds. In CIFAR-10-C, ShiftEx completes adaptation in 9–19 rounds, while baselines remain stagnant beyond 51 rounds. On FEMNIST and FashionMNIST, ShiftEx achieves recovery within 0–42 rounds, including several windows where adaptation completes in under 5 rounds—markedly outperforming methods that require 50+ rounds without convergence. These trends are also evident in the convergence curves shown in Figures 3(a) through 4(b), where ShiftEx consistently demonstrates steeper accuracy recovery slopes and faster stabilization.
@@
Interestingly, OORT shows a smaller initial accuracy drop, but this reflects underreaction rather than robustness to shift. Its static party selection masks distribution shifts, leading to limited recovery and lower final accuracy. In contrast, ShiftEx detects shifts early and creates and trains new experts, causing a brief dip but enabling rapid recovery and higher long-term performance. These results confirm that ShiftEx does not trade off speed for accuracy—it excels in both.
@@ 8. Related Work
Continual and Temporal Adaptation in FL. To go beyond static training assumptions, recent efforts focus on temporal and streaming FL. Adaptive FL with mixture models (Chen et al., 2024) and time-aware multi-branch networks (Zhu et al., 2021) attempt to model evolving distributions, but rely on fixed priors. Streaming FL techniques (Marfoq et al., 2023; Wang et al., 2023) leverage online data collection and cache update rules to manage distributional drift. Other approaches target evolving concepts through semi-supervised prototypes (Mawuli et al., 2023), drift-aware reweighting (Casado et al., 2022), and dynamic party clustering (Jothimurugesan et al., 2023; Li et al., 2024). Federated continual learning (FCL) (Hamedi et al., 2025; Wang et al., 2024) and heterogeneity-aware continual learning (Criado et al., 2022) consolidate strategies for learning under drift while preserving stability and privacy. Lightweight two-stream architectures (Yao et al., 2018) further reduce communication under shift. Distinct from these, ShiftEx offers a modular MoE framework that continuously adapts to both abrupt and gradual changes without relying on explicit drift signals or fixed temporal structures.
@@ 9. Conclusion
We present a shift-aware mixture of experts framework for federated learning under non-stationary data, addressing covariate and label shifts in streaming environments. Through shift detection, latent memory for expert reuse, and facility location-based optimization, our method enables dynamic expert instantiation and scalable adaptation. Extensive experiments show significant gains in accuracy and adaptation speed over FL baselines, offering a practical, privacy-preserving middleware for real-world deployments. Traditional FL middleware approaches assume relatively stable workload characteristics, but our framework addresses the unique challenge of building middleware that can detect, respond to, and manage distributional shifts in real-time across federated participants. The system implements key middleware patterns, including event-driven architectures (responding to detected shifts), load balancing (distributing clients across experts), and service discovery (expert reuse and consolidation). The windowing-based stream processing and latent memory mechanisms function as middleware services that enable scalable, real-time adaptation without requiring application-level awareness of underlying distributional changes. Limitations include potential expert pool growth under extreme heterogeneity, reliance on frozen encoders, and honest client reporting. Future work will explore expert compression via online distillation, secure reporting under adversarial settings, and tighter integration with hardware enclaves for end-to-end privacy.
Acknowledgements.

### v2

@@ 7. Results and Discussion
We present the numerical results for Adaptation Time, Accuracy Drop, and Max Accuracy in Tables 1 and 2. We report standard deviations for Accuracy Drop and Max Accuracy across six runs. For Adaptation Time, standard deviations are omitted since the variation is negligible, generally amounting to only a few rounds. To visualize convergence behavior, we plot Accuracy (%) over communication rounds in Figures 3(a) through 4(b). Figures 5(a) through 6(b) illustrate the peak accuracy achieved by each method in every window, where error bars represent the variability (standard deviation) observed across six runs. Finally, to illustrate how ShiftEx dynamically adapts to distributional changes, we visualize expert assignment distributions across windows in Figures 7(a) through 8(b). These plots show how ShiftEx begins with a single expert in W0 and progressively creates and trains new experts as shifts are detected, enabling specialized adaptation to new data regimes.
ShiftEx reduces adaptation time under distribution shifts. Across all datasets, ShiftEx achieves significantly faster adaptation than all baselines. On the large-scale FMoW dataset (Table 1), where most baselines fail to recover within 51 rounds, ShiftEx consistently recovers within 1–13 rounds across all windows. In TinyImagenet-C (Table 2), recovery time drops to 27–29 rounds, whereas other methods plateau with no recovery even after 40 rounds. In CIFAR-10-C, ShiftEx completes adaptation in 9–19 rounds, while baselines remain stagnant beyond 51 rounds. On FEMNIST and FashionMNIST, ShiftEx achieves recovery within 0–42 rounds, including several windows where adaptation completes in under 5 rounds—markedly outperforming methods that require 50+ rounds without convergence. These trends are also evident in the convergence curves shown in Figures 3(a) through 4(b), where ShiftEx consistently demonstrates steeper accuracy recovery slopes and faster stabilization. Even in scenarios where all clients simultaneously experience a shift, clustering still groups similar clients together and triggers creation of new experts, ensuring robustness without collapse.
@@
Behavior of baselines: Figures 3(a) through 4(b) illustrate accuracy trends across different baselines, and these behaviors can be understood in relation to the types of distribution shifts. FedProx and Fielding rely on a single global model or static label-aware clustering and therefore struggle under covariate shifts, leading to sharp drops and limited recovery. OORT shows a smaller initial accuracy drop, but this reflects underreaction rather than robustness: its utility-based client selection masks distributional changes, preventing effective adaptation and resulting in limited recovery and lower final accuracy. FedDrift performs better under moderate covariate drift, as seen with controlled corrections on CIFAR-10-C, since it clusters clients by loss patterns, but it lacks explicit handling of label shifts and cannot adapt when class distributions skew significantly with the other large-scale datasets like FMoW and Tiny-ImageNet-C. In contrast, ShiftEx explicitly detects both covariate and label shifts and instantiates new experts when needed, which explains its accuracy dips followed by rapid recovery and consistently higher post-shift accuracy.
@@ 8. Related Work
Continual and Temporal Adaptation in FL. To go beyond static training assumptions, recent efforts focus on temporal and streaming FL. Adaptive FL with mixture models (Chen et al., 2024) and time-aware multi-branch networks (Zhu et al., 2021) attempt to model evolving distributions, but rely on fixed priors. Streaming FL techniques (Marfoq et al., 2023; Wang et al., 2023a) leverage online data collection and cache update rules to manage distributional drift. Other approaches target evolving concepts through semi-supervised prototypes (Mawuli et al., 2023), drift-aware reweighting (Casado et al., 2022), and dynamic party clustering (Jothimurugesan et al., 2023; Li et al., 2024). Federated continual learning (Hamedi et al., 2025; Wang et al., 2024) and heterogeneity-aware continual learning (Criado et al., 2022a) address drift while preserving stability and privacy, while lightweight two-stream architectures (Yao et al., 2018) cut communication under shift. Distinct from these, ShiftEx offers a modular MoE framework that continuously adapts to both abrupt and gradual changes without relying on explicit drift signals or fixed temporal structures.
Federated continual learning (FCL). FCL extends FL to dynamic, non-stationary environments by enabling models to adapt to sequential tasks or streaming data while mitigating catastrophic forgetting. Early work such as FedWeIT (Yoon et al., 2021a) proposed task-adaptive parameter decomposition and selective inter-client transfer, while FedCIL (Dong et al., 2022) introduced prototype-based distillation for class-incremental FL. Rehearsal and regularization approaches, including FedPMR (Wang et al., 2023b), preserve prior knowledge through probability alignment and parameter-consistency constraints. More recently, drift-aware device management for IoT-based FL (Zhou et al., 2024) integrates server-side drift detection with exemplar-based continual learning to balance resource constraints with adaptation. Other federated continual and drift-robust methods include IFCA, which clusters clients into static groups (Ghosh et al., 2020), and test-time adaptation approaches such as FedTHE (Jiang and Lin, 2023), ATP (Bao et al., 2023), and FedCTTA (Hasan Rajib et al., 2025). Surveys further highlight FCL’s dual challenge of preserving prior knowledge and fusing heterogeneous client experiences (Yang et al., 2024; Criado et al., 2022b). In contrast, ShiftEx contributes a MoE middleware that performs distributional client routing without input-level gating, constrains expert growth via expert consolidation, and narrows the leakage surface by operating only on aggregate latent statistics inside a TEE.
@@ 9. Conclusion
We present a shift-aware mixture of experts framework for federated learning under non-stationary data, addressing covariate and label shifts in streaming environments. Through shift detection, latent memory for expert reuse, and facility location-based optimization, our method enables dynamic expert instantiation and scalable adaptation. Extensive experiments show significant gains in accuracy and adaptation speed over FL baselines, offering a practical, privacy-preserving middleware for real-world deployments. Traditional FL middleware approaches assume relatively stable workload characteristics, but our framework addresses the unique challenge of building middleware that can detect, respond to, and manage distributional shifts in real-time across federated participants. The system implements key middleware patterns, including event-driven architectures (responding to detected shifts), load balancing (distributing clients across experts), and service discovery (expert reuse and consolidation). The windowing-based stream processing and latent memory mechanisms function as middleware services that enable scalable, real-time adaptation without requiring application-level awareness of underlying distributional changes. Limitations include potential expert pool growth under extreme heterogeneity, reliance on frozen encoders, and honest client reporting. Under extreme heterogeneity, the number of experts may proliferate; however, our consolidation step periodically merges similar experts to mitigate this risk. Future work will explore expert compression via online distillation, secure reporting under adversarial settings, and tighter integration with hardware enclaves for end-to-end privacy.